### v1

@@ 3.1 Overview of the algorithm for identifying fibers and bases
Following the approach developed in Huang-WT-fibers, for each point $p\in L$, we define the function $s(p)=\text{max}_{v\in V}\langle p,v\rangle$ and define sets $S_{i}=\{p\in L:s(p)=i\}$. Then, $\nabla$ contains a 2D reflexive subpolytope if and only if it satisfies one of the following conditions:
@@
Table 3: The sizes of sets $|S_{i}|$ used to identify each fiber Huang-WT-fibers.
@@
In all cases, the resulting sequence of intersection numbers corresponds precisely to one of the Morrison-Taylor (MT) bases identified in MorrisonTaylorToric.
@@ 3.2 Identifying equivalence of fibrations under polytope automorphisms
Fix a 4D reflexive polytope $\nabla$. For fibrations with a smooth base, the data defining a fibration structure are simply the fiber type, in the conventions of Table 2, and the base type, i.e. the member of the MT list MorrisonTaylorToric with the same toric intersection structure as the base of the fibration. For fibrations with singular bases, in addition to the topological type of the resolved base, we also include a classification of the curves which are blown down to yield the singular base. Thus, in total, a fibration structure is identified by five pieces of information: the ambient polytope, the fiber type, the base type, whether the base is smooth, and a (possibly empty) list of curves that have been blown down. [^5] An additional complication arises when the base of the fibration is singular and the resolved base itself has nontrivial automorphisms, as then the specification of blowdowns becomes ambiguous. In this rare occurrence, we resolve this ambiguity by selecting the lexicographically minimal representative of the specification of blown down curves under the base automorphisms.
Geometrically, automorphism-equivalent fibrations should be identified. Furthermore, from the physics perspective, as emphasized in Braun:2011ux, because compactifying F-theory on elliptic fibrations defined by automorphism equivalent subpolytopes gives identical physical theories, it is also natural in this context to count the number of automorphism equivalence classes of fibrations, rather than the total number of fibrations. Thus, given a 4D reflexive subpolytope $\nabla$, once we have identified the fibers and bases of all of its elliptic fibrations using the algorithm above, we also compute the automorphisms of $\nabla$ and check whether any subpolytopes are mapped into each other by those automorphisms. In practice, we first sort the subpolytopes into fibration structures, and only compute automorphisms if at least one fibration structure has more than one representative.
@@ 3.3 Computational details
In the last two subsections, we have spelled out an algorithm for completely classifying the automorphism-inequivalent toric elliptic fibrations of a 4D reflexive polytope $\nabla$ from the Kreuzer-Skarke list. We have applied this algorithm, which we have implemented in both CYTools Demirtas:2022hqf and Julia bezanson2017julia, to all 473,800,776 4D reflexive polytopes to obtain a complete classification of all toric elliptic fibrations. Before presenting the results of our analysis in §4, we briefly outline the technical details of this computation, which was carried out on MIT’s subMIT cluster Bendavid:2025arv.
The KS database KS classifies polytopes by the number of vertices of their dual polytope. Thus, the file v05.gz contains the 1,561 polytopes whose duals have five vertices each. [^6] One immediate way to verify that the KS database organizes its files by the number of vertices of the dual, rather than the number of vertices of the polytope itself, is to query it for polytopes with $h^{1,1}=1,h^{2,1}=101$; doing so returns the vertices of the mirror quintic rather than those of the quintic itself. We have kept this organizational scheme throughout our analysis, including in the database accompanying this publication, but have split the vnn.gz files into more manageable files of (up to) one million polytopes each, labeled vnn-1.txt and so on, when appropriate.
@@ 3.4 Organization of results and nomenclature
The results of the analysis are organized in a set of data files that are publicly available at archive. The details of the formatting of the files are described in more detail in the README.txt file associated with the archive, but we give a broad description here.
In the remainder of this paper and the Zenodo database, we refer to a given polytope using the syntax “vxxxx-yyyyyy,” where (following the format used in the KS database KS), $xx$ denotes the number of vertices of the dual polytope (2 digits, ranging from 05–36) and yyyyyy denotes indexing in the order of the KS database for polytopes with that number of vertices (varying numbers of digits). Fibers 1–16 are ordered as in Klevers:2014bqa and Figure 1, and bases 1–61,539 are ordered canonically as in the database associated with MorrisonTaylorToric, i.e. they are ordered first by $h^{1,1}(B_{2})$, and entries at each value of $h^{1,1}$ are ordered canonically according to (the negatives of the) intersection numbers of the toric base divisors, so e.g. bases $1,2,\ldots,13$ correspond to $\mathbb{P}^{2},\mathbb{F}_{0},\ldots,\mathbb{F}_{12}$). For convenience, that database is also included in the Zenodo archive.
@@
[base singular?, base ID, $h^{1,1}(B)$, [blown down curves],...
@@
   [[base 1 singular?, base 1 ID, $h^{1,1}$(base 1), [blown down curves in base 1],
@@
   [base 2 singular?, base 2 ID, $h^{1,1}$(base 2), [blown down curves in base 2],
@@
    $F_{10}$ fibered over base 150, which has $h^{1,1}(B)=6$, with no blowdowns. There are two fibrations with this structure, and they are related by an automorphism.
@@
    $F_{1}$ fibered over base 4162, which has $h^{1,1}(B)=14$, with blowdowns. The (negative) intersection form of this base is
@@ 4.1 Fibrations per polytope
The average number of fibrations per polytope is 4.78, of which 4.75 are automorphism-inequivalent. It is interesting to compare this with the average number of “obvious” fibrations for CICYs, which was found to be 9.85 in Anderson:2016cdu.
@@ 4.2 Fiber types
We find fibrations with each of the sixteen fibers. This is guaranteed since for any pair of fibers $F,F^{\prime}$ we can construct a reflexive 4D polytope as a product (with rays $(v\in F,v^{\prime}\in F^{\prime})$) or a “sum” (with rays $(v\in F,0),(0,v^{\prime}\in F^{\prime})$) of these two fibers.  [^8] A detailed analysis of these polytopes and their behavior under mirror symmetry will be given in ot-forthcoming. The distribution of fiber types is given in Table 4 and illustrated graphically in Figure 3. We find that $F_{6}$ is the most common fiber ($\sim 18\%$), with 403,991,733 automorphism-inequivalent fibrations, while $F_{16}$ is the least common fiber ($\sim 0.01\%$), with only 182,915 inequivalent fibrations. On the other hand, for polytopes with large $h^{1,1}(X)$, $F_{10}$ dominates, as can be seen in Figure 4; the highest $h^{1,1}$ of a polytope that has no $F_{10}$ fibers is $h^{1,1}=227$. The dominance of fiber $F_{6}$ was to us somewhat unexpected; because every base admits at least one fibration with fiber $F_{10}$ (the standard stacking, §2.4), and this fiber is closest to the generic Weierstrass model, one might have expected $F_{10}$ to dominate. Indeed, much of the previous analysis of fibrations focused implicitly on $F_{10}$ fibers. We explore the geometry of fibrations with fiber $F_{6}$ in further detail in §5.4.
@@ 4.3 Bases
We find that each of the 61,539 toric bases on the MT list appears as the base for at least one fibration; 625 bases appear exactly once. This is unsurprising; as explained in §2.4, one can construct the generic elliptic fibration over any toric base by taking (the dual of the dual of) the standard stacking of $F_{10}$ over that base.
While all bases appear, there is a marked statistical preference towards bases with small $h^{1,1}(B)$. As discussed in more detail in §5.4.1, this is likely because bases with larger $h^{1,1}(B)$ have a large number of non-Higgsable clusters that strongly constrain the possibility of additional tuned gauge factors. The most common base, which is entry 72 on the MT list, is a generalized del Pezzo surface $gdP_{4}$, with $h^{1,1}(B)=5$, and is the base for 5.1% (114,961,701 total, 114,019,846 inequivalent) of all fibrations; this is a weak Fano base, and is the toric variety defined by the $F_{12}$ polytope, as in Figure 1. Plotting the number of fibrations over a base as a function of $h^{1,1}(B)$, as shown in Figure 5, we find a distinct peak at small $h^{1,1}(B)$. In particular, 82.5% (1,868,680,827 total, 1,857,694,025 inequivalent) of fibrations have $h^{1,1}(B)\leq 10$, and 99.4% (2,251,269,536 total, 2,237,134,428 inequivalent) have $h^{1,1}(B)\leq 25$. We have plotted the Hodge numbers of all elliptic fibrations over a number of specific bases with small $h^{1,1}(B)$ in Figure 6.
Figure 5: The mean, minimum, maximum and $1\sigma$ deviation of the number of inequivalent fibrations over a base $B$ with a given $h^{1,1}(B)$ are plotted. Note that there are more structures fibered over bases with small $h^{1,1}(B).$
@@
Moreover, as we see in Fig. 7, the only fiber type that appears fibered over a base with $h^{1,1}(B_{2})>100$ is $F_{10}$. This corroborates the expectation (c.f. §2.4) that for bases with large $h^{1,1}(B_{2})$, the only available elliptic fibration is the generic elliptic fibration, associated with the standard stacking of the fiber $F_{10}$.
Figure 7: The number of inequivalent fibrations of each fiber type as a function of $h^{1,1}(B)$. Note that at large $h^{1,1}(B)$, only fibrations with fiber $F_{10}$ are seen.
@@ 4.3.1 Weak Fano and non-weak-Fano bases
As discussed in §2.4, of the 61,539 toric bases, 16 are weak Fano; these exactly correspond to the 16 toric fibers. Some summary statistics of fibrations with weak Fano bases are given in Table 5. We find that fibrations with weak Fano bases are disproportionately common, comprising 26% (588,457,836 total, 584,970,461 inequivalent) of all fibrations. In particular, the four most common bases are all weak Fano. Intuitively, this makes sense; weak Fano bases lack NHCs, and so one has additional freedom to tune gauge groups over weak Fano bases relative to non-weak-Fano bases. We show the distribution of fiber types in fibrations over the sixteen weak Fano bases in Figure 8(a).
@@ 73 $F_{11}$ 5 82,902,723 87,426,934 87,830,312
Table 6: Summary statistics for fibrations with the sixteen most common non-weak-Fano bases. We specify the bases by their labels in the MT list. For each base, we have indicated the Hodge number $h^{1,1}(B)$ and the (negative of) its intersection form, as well as the number of fibration structures, the number of inequivalent fibrations, and the total number of fibrations with that base. Note that the first five most common bases in the list differ from weak Fano by containing only a single curve of self-intersection -3 or -4.
@@ 4.3.2 Singular Bases
Of the 61,539 toric bases, only 17,961 appear as the resolution of a singular base in a genus-one fibration. The most common base is base 417, a non-weak-Fano base with $h^{1,1}(B)=8$ and which appears as (the resolution of) the base in 3.3% (372,445 total, 370,503 inequivalent) of such fibrations. The most common overall base, MT base 72 or equivalently the fiber $F_{12}$, appears as a singular base in only 1% (116,300 total, 115,138 inequivalent) of such fibrations. In spite of this shift, the statistical preference for small $h^{1,1}(B)$ remains; the largest singular base has $h^{1,1}(B)=92$, and 97.7% (10,929,212 total, 10,882,539) of such fibrations have $h^{1,1}(B)\leq 25$.
@@ 4.4 Rank increase
$\delta_{\text{tuned}}:=h^{1,1}(X)-\left[h^{1,1}(B)+\operatorname{rk}\left(\text{NHC}\right)\right]-1.$ (33)
@@
While for a weak Fano base these two notions coincide, for a non-weak-Fano base the tuned rank increase more closely quantifies how much tuning the fibration has by subtracting out a universal contribution from the NHC; we therefore find it to be a more useful definition. We have computed both the total and tuned rank increase for each fibration, but for the remainder of this section we will comment only on the tuned rank increases; for brevity, we will refer to the tuned rank increase simply as the rank increase in what follows.
Note that, for elliptic fibrations, the rank increase is nonnegative by Shioda-Tate Wazir. Nevertheless, we find fibrations with negative rank increase; this occurs only in genus-one fibrations, and is relatively rare, occurring in a total of only 3,868,168 fibrations (3,829,298 inequivalent). The appearance of negative rank increase reflects the fact that the Hodge numbers of the Jacobian fibration of a genus-one fibration are generally not equal to those of the genus-one fibered threefold Anderson:2023wkr; Anderson:2023tfy. To avoid this complication, in what follows we will consider only genuine elliptic fibrations.
@@
Figure 9: The Hodge numbers of the generic elliptic fibration over each base in the the MT list. For each Hodge pair, $h^{1,1}(B)$ is indicated by the color. 7,524 of the 30,108 unique Hodge pairs in the Kreuzer-Skarke list are realized in this way.
The remaining 1,828,049,985 elliptic fibrations (1,817,131,593 inequivalent) have positive rank increase. We have plotted the number of inequivalent elliptic fibrations of each rank increase in Figure 10. The largest bases that appear with positive rank increase are bases 61,521 and 61,523, each of which has $h^{1,1}(B)=150$ and appears as the base for a single fibration with $\delta_{\text{tuned}}=1$. The most common rank increase is $\delta_{\text{tuned}}=15$, which occurs in 101,205,632 elliptic fibrations (100,589,846 inequivalent). We find that large rank increases are extremely rare; only 3,933 (3,931 inequivalent) fibrations have $\delta_{\text{tuned}}\geq 200$.
@@
We have plotted the relative density of each of the thirteen fiber types of elliptic fibrations in Figure 11; from this figure, it is clear that fibrations by fiber $F_{13}$ dominate at large $\delta_{r}$. The largest rank increase is $\delta_{r}=463$, found in an $F_{13}$ fibration over base 680 in polytope v05-393, the unique polytope with $h^{1,1}(X)=491$; we discuss this polytope and its fibrations in §5.6. Indeed, this fiber-base pair is universal at large $\delta_{r}$; the largest rank increase that does not come from an $F_{13}$ fibration over base 680 is 350, found in an $F_{13}$ fibration over base 679 in a polytope v06-2907, which has $h^{1,1}(10)=376$, $h^{2,1}(X)=10$. By Eq. 14c, given an elliptic fibration in a CY3 with Hodge numbers $h^{1,1},h^{2,1}$, we can sometimes trade one unit of $h^{1,1}$ for 29 units of $h^{2,1}$. The prevalence of fibrations of $F_{13}$ over base 680 at large $\delta_{r}$ can (at least in part) be explained by the mirror of this process (which will also be described further in ot-forthcoming): their Hodge numbers are given by $h^{1,1}(X)=491-29n$, $h^{2,1}(X)=11+n$, as expected for the images of the $F_{13}$ fibration in the $h^{1,1}=491$ polytope under the mirror of these $\Delta{T}=1$ transitions. Indeed, in Figure 6 it is clear that fibrations over base 680 dominate the large $h^{1,1}(X)$ regime.
Finally, we note a curious fact: each polytope hosting one of these large rank-increase fibrations of fiber $F_{13}$ over base 680 also possesses a second fibration, namely the generic elliptic fibration over a base with very large $h^{1,1}(B)$. For instance, as discussed in §5.6, the polytope with $h^{1,1}(X)=491$ also hosts the generic elliptic fibration over the largest toric base. Although we currently cannot explain this empirical phenomenon, we hope that the ideas of Anderson:2016cdu can be applied to these pairs of fibrations.

### v2

@@ 3.1 Overview of the algorithm for identifying fibers and bases
Following the approach developed in Huang and Taylor (2020), for each point $p\in L$, we define the function $s(p)=\text{max}_{v\in V}\langle p,v\rangle$ and define sets $S_{i}=\{p\in L:s(p)=i\}$. Then, $\nabla$ contains a 2D reflexive subpolytope if and only if it satisfies one of the following conditions:
@@
Table 3: The sizes of sets $|S_{i}|$ used to identify each fiber Huang and Taylor (2020).
@@
In all cases, the resulting sequence of intersection numbers corresponds precisely to one of the Morrison-Taylor (MT) bases identified in Morrison and Taylor (2012c).
This algorithm gives the maximal toric base for an elliptic CY3 associated with a triangulation of $\nabla$. In many cases, there are also triangulations corresponding to flop phases in which fewer base divisors are compatible with the toric morphism, and there is thus a non-flat fibration to a base with smaller $h^{1,1}(B)$. [^6] For example, the polytope associated with the standard stacking over Hirzebruch $\mathbb{F}_{1}$ has a triangulation giving the generic elliptic CY3 over $\mathbb{F}_{1}$, but there is also a triangulation giving a non-flat fibration over $\mathbb{P}^{2}$ Nally and Taylor . In general, we find it useful to organize fibrations using the phase where the base has maximal $h^{1,1}(B)$; this is also part of the reason that it is useful to track $h^{1,1}_{*}(B)$, as defined in (4), which extends this consideration to flop phases where non-flat fibers automatically forced from the structure of the base are traded for base divisors, but the base is no longer toric.
@@ 3.2 Identifying equivalence of fibrations under polytope automorphisms
Fix a 4D reflexive polytope $\nabla$. For fibrations with a smooth base, the data defining a fibration structure are simply the fiber type, in the conventions of Table 2, and the base type, i.e. the member of the MT list Morrison and Taylor (2012c) with the same toric intersection structure as the base of the fibration. For fibrations with singular bases, in addition to the topological type of the resolved base, we also include a classification of the curves which are blown down to yield the singular base. Thus, in total, a fibration structure is identified by five pieces of information: the ambient polytope, the fiber type, the base type, whether the base is smooth, and a (possibly empty) list of curves that have been blown down. [^7] An additional complication arises when the base of the fibration is singular and the resolved base itself has nontrivial automorphisms, as then the specification of blowdowns becomes ambiguous. In this rare occurrence, we resolve this ambiguity by selecting the lexicographically minimal representative of the specification of blown down curves under the base automorphisms.
Geometrically, automorphism-equivalent fibrations should be identified. Furthermore, from the physics perspective, as emphasized in Braun (2013), because compactifying F-theory on elliptic fibrations defined by automorphism equivalent subpolytopes gives identical physical theories, it is also natural in this context to count the number of automorphism equivalence classes of fibrations, rather than the total number of fibrations. Thus, given a 4D reflexive subpolytope $\nabla$, once we have identified the fibers and bases of all of its elliptic fibrations using the algorithm above, we also compute the automorphisms of $\nabla$ and check whether any subpolytopes are mapped into each other by those automorphisms. In practice, we first sort the subpolytopes into fibration structures, and only compute automorphisms if at least one fibration structure has more than one representative.
@@ 3.3 Computational details
In the last two subsections, we have spelled out an algorithm for completely classifying the automorphism-inequivalent toric elliptic fibrations of a 4D reflexive polytope $\nabla$ from the Kreuzer-Skarke list. We have applied this algorithm, which we have implemented in both CYTools Demirtas et al. (2022) and Julia Bezanson et al. (2017), to all 473,800,776 4D reflexive polytopes to obtain a complete classification of all toric elliptic fibrations. Before presenting the results of our analysis in §4, we briefly outline the technical details of this computation, which was carried out on MIT’s subMIT cluster Bendavid and others (2025).
The KS database "Kreuzer and Skarke classifies polytopes by the number of vertices of their dual polytope. Thus, the file v05.gz contains the 1,561 polytopes whose duals have five vertices each. [^8] One immediate way to verify that the KS database organizes its files by the number of vertices of the dual, rather than the number of vertices of the polytope itself, is to query it for polytopes with $h^{1,1}=1,h^{2,1}=101$; doing so returns the vertices of the mirror quintic rather than those of the quintic itself. We have kept this organizational scheme throughout our analysis, including in the database accompanying this publication, but have split the vnn.gz files into more manageable files of (up to) one million polytopes each, labeled vnn-1.txt and so on, when appropriate.
@@ 3.4 Organization of results and nomenclature
The results of the analysis are organized in a set of data files that are publicly available at Abbasi et al. (2025). The details of the formatting of the files are described in more detail in the README.txt file associated with the archive, but we give a broad description here.
In the remainder of this paper and the Zenodo database, we refer to a given polytope using the syntax “vxxxx-yyyyyy,” where (following the format used in the KS database "Kreuzer and Skarke ), $xx$ denotes the number of vertices of the dual polytope (2 digits, ranging from 05–36) and yyyyyy denotes indexing in the order of the KS database for polytopes with that number of vertices (varying numbers of digits). Fibers 1–16 are ordered as in Klevers et al. (2015) and Figure 1, and bases 1–61,539 are ordered canonically as in the database associated with Morrison and Taylor (2012c), i.e. they are ordered first by $h^{1,1}(B_{2})$, and entries at each value of $h^{1,1}$ are ordered canonically according to (the negatives of the) intersection numbers of the toric base divisors, so e.g. bases $1,2,\ldots,13$ correspond to $\mathbb{P}^{2},\mathbb{F}_{0},\ldots,\mathbb{F}_{12}$). For convenience, that database is also included in the Zenodo archive.
@@
[base singular?, base ID, $h_{*}^{1,1}(B)$, [blown down curves],...
where $h^{1,1}_{*}(B)$ is given by (4).
@@
   [[base 1 singular?, base 1 ID, $h_{*}^{1,1}$(base 1), [blown down curves in base 1],
@@
   [base 2 singular?, base 2 ID, $h_{*}^{1,1}$(base 2), [blown down curves in base 2],
@@
    $F_{10}$ fibered over base 150, which has $h^{1,1}_{*}(B)=h^{1,1}(B)=6$, with no blowdowns. There are two fibrations with this structure, and they are related by an automorphism.
@@
    $F_{1}$ fibered over base 4162, which has $h^{1,1}_{*}(B)=h^{1,1}(B)=14$, with blowdowns. The (negative) intersection form of this base is
@@ 4.1 Fibrations per polytope
The average number of fibrations per polytope is 4.78, of which 4.75 are automorphism-inequivalent. It is interesting to compare this with the average number of “obvious” fibrations for CICYs, which was found to be 9.85 in Anderson et al. (2016a).
@@ 4.2 Fiber types
We find fibrations with each of the sixteen fibers. This is guaranteed since for any pair of fibers $F,F^{\prime}$ we can construct a reflexive 4D polytope as a product (with rays $(v\in F,v^{\prime}\in F^{\prime})$) or a “sum” (with rays $(v\in F,0),(0,v^{\prime}\in F^{\prime})$) of these two fibers.  [^10] A detailed analysis of these polytopes and their behavior under mirror symmetry will be given in Oehlmann and Taylor . The distribution of fiber types is given in Table 4 and illustrated graphically in Figure 3. We find that $F_{6}$ is the most common fiber ($\sim 18\%$), with 403,991,733 automorphism-inequivalent fibrations, while $F_{16}$ is the least common fiber ($\sim 0.01\%$), with only 182,915 inequivalent fibrations. On the other hand, for polytopes with large $h^{1,1}(X)$, $F_{10}$ dominates, as can be seen in Figure 4; the highest $h^{1,1}$ of a polytope that has no $F_{10}$ fibers is $h^{1,1}=227$. The dominance of fiber $F_{6}$ was to us somewhat unexpected; because every base admits at least one fibration with fiber $F_{10}$ (the standard stacking, §2.4), and this fiber is closest to the generic Weierstrass model, one might have expected $F_{10}$ to dominate. Indeed, much of the previous analysis of fibrations focused implicitly on $F_{10}$ fibers. We explore the geometry of fibrations with fiber $F_{6}$ in further detail in §5.4.
@@ 4.3 Bases
In this section we discuss the statistics of fibrations, in particular how fibrations are distributed as a function of the effective Hodge number of the base $h^{1,1}_{*}(B)$ defined in (4). We find that each of the 61,539 toric bases on the MT list appears as the base for at least one fibration; 625 bases appear exactly once. This is unsurprising; as explained in §2.4, one can construct the generic elliptic fibration over any toric base by taking (the dual of the dual of) the standard stacking of $F_{10}$ over that base.
While all bases appear, there is a marked statistical preference towards bases with small $h^{1,1}_{*}(B)$. As discussed in more detail in §5.4.1, this is likely because bases with larger $h^{1,1}_{*}(B)$ have a large number of non-Higgsable clusters that strongly constrain the possibility of additional tuned gauge factors. The most common base, which is entry 72 on the MT list, is a generalized del Pezzo surface $gdP_{4}$, with $h^{1,1}_{*}(B)=h^{1,1}(B)=5$, and is the base for 5.1% (114,961,701 total, 114,019,846 inequivalent) of all fibrations; this is a weak Fano base, and is the toric variety defined by the $F_{12}$ polytope, as in Figure 1. Plotting the number of fibrations over a base as a function of $h_{*}^{1,1}(B)$, as shown in Figure 5, we find a distinct peak at small $h^{1,1}_{*}(B)$. In particular, 82.5% (1,868,680,827 total, 1,857,694,025 inequivalent) of fibrations have $h^{1,1}_{*}(B)\leq 10$, and 99.4% (2,251,269,536 total, 2,237,134,428 inequivalent) have $h^{1,1}_{*}(B)\leq 25$. We have plotted the Hodge numbers of all elliptic fibrations over a number of specific bases with small $h^{1,1}_{*}(B)$ in Figure 6.
Figure 5: The mean, minimum, maximum and $1\sigma$ deviation of the number of inequivalent fibrations over a base $B$ with a given $h^{1,1}_{*}(B)$ are plotted. Note that there are more structures fibered over bases with small $h^{1,1}_{*}(B).$
@@
Moreover, as we see in Fig. 7, the only fiber type that appears fibered over a base with $h^{1,1}_{*}(B_{2})>100$ is $F_{10}$. This corroborates the expectation (c.f. §2.4) that for bases with large $h^{1,1}_{*}(B_{2})$, the only available elliptic fibration is the generic elliptic fibration, associated with the standard stacking of the fiber $F_{10}$.
Figure 7: The number of inequivalent fibrations of each fiber type as a function of $h^{1,1}_{*}(B)$. Note that at large $h^{1,1}_{*}(B)$, only fibrations with fiber $F_{10}$ are seen.
@@ 4.3.1 Weak Fano and non-weak-Fano bases
As discussed in §2.4, of the 61,539 toric bases, 16 are weak Fano; these exactly correspond to the 16 toric fibers. Some summary statistics of fibrations with weak Fano bases are given in Table 5. We find that fibrations with weak Fano bases are disproportionately common, comprising 26% (588,457,836 total, 584,970,461 inequivalent) of all fibrations. In particular, the four most common bases are all weak Fano. Intuitively, this makes sense; weak Fano bases lack NHCs, and so one has additional freedom to tune gauge groups over weak Fano bases relative to non-weak-Fano bases. We show the distribution of fiber types in fibrations over the sixteen weak Fano bases in Figure 8(a). Note that for weak Fano bases $h^{1,1}(B)=h^{1,1}_{*}(B)$ since there are no toric curves of self-intersection less than -2.
@@ 73 $F_{11}$ 5 82,902,723 87,426,934 87,830,312
Table 6: Summary statistics for fibrations with the sixteen most common non-weak-Fano bases. We specify the bases by their labels in the MT list. For each base, we have indicated the Hodge number $h^{1,1}(B)=h^{1,1}_{*}(B)$ and the (negative of) its intersection form, as well as the number of fibration structures, the number of inequivalent fibrations, and the total number of fibrations with that base. Note that the first five most common bases in the list differ from weak Fano by containing only a single curve of self-intersection -3 or -4.
@@ 4.3.2 Singular Bases
Of the 61,539 toric bases, only 17,961 appear as the resolution of a singular base in a genus-one fibration. The most common base is base 417, a non-weak-Fano base with $h^{1,1}_{*}(B)=h^{1,1}(B)=8$ and which appears as (the resolution of) the base in 3.3% (372,445 total, 370,503 inequivalent) of such fibrations. The most common overall base, MT base 72 or equivalently the fiber $F_{12}$, appears as a singular base in only 1% (116,300 total, 115,138 inequivalent) of such fibrations. In spite of this shift, the statistical preference for small $h^{1,1}_{*}(B)$ remains; the largest singular base has $h^{1,1}_{*}(B)=92$, and 97.7% (10,929,212 total, 10,882,539) of such fibrations have $h^{1,1}_{*}(B)\leq 25$.
@@ 4.4 Rank increase
$\delta_{\text{tuned}}:=h^{1,1}(X)-\left[h^{1,1}_{*}(B)+\operatorname{rk}\left(\text{NHC}\right)\right]-1.$ (34)
@@
While for a weak Fano base these two notions coincide, for a non-weak-Fano base the tuned rank increase more closely quantifies how much tuning the fibration has by subtracting out a universal contribution from the NHC (including both contributions from non-Higgsable gauge groups and non-Higgsable SCFT’s); we therefore find it to be a more useful definition. We have computed both the total and tuned rank increase for each fibration, but for the remainder of this section we will comment only on the tuned rank increases; for brevity, we will refer to the tuned rank increase simply as the rank increase in what follows.
Note that, for elliptic fibrations, the rank increase is nonnegative by Shioda-Tate Wazir. Nevertheless, we find fibrations with negative rank increase; this occurs only in genus-one fibrations, and is relatively rare, occurring in a total of only 3,868,168 fibrations (3,829,298 inequivalent). The appearance of negative rank increase reflects the fact that the Hodge numbers of the Jacobian fibration of a genus-one fibration are generally not equal to those of the genus-one fibered threefold Anderson et al. (2024, 2023). To avoid this complication, in what follows we will consider only genuine elliptic fibrations.
@@
Figure 9: The Hodge numbers of the generic elliptic fibration over each base in the the MT list. For each Hodge pair, $h^{1,1}_{*}(B)$ is indicated by the color. 7,524 of the 30,108 unique Hodge pairs in the Kreuzer-Skarke list are realized in this way.
The remaining 1,828,049,985 elliptic fibrations (1,817,131,593 inequivalent) have positive rank increase. We have plotted the number of inequivalent elliptic fibrations of each rank increase in Figure 10. The largest bases that appear with positive rank increase are bases 61,521 and 61,523, each of which has $h^{1,1}_{*}(B)=150$ and appears as the base for a single fibration with $\delta_{\text{tuned}}=1$. The most common rank increase is $\delta_{\text{tuned}}=15$, which occurs in 101,205,632 elliptic fibrations (100,589,846 inequivalent). We find that large rank increases are extremely rare; only 3,933 (3,931 inequivalent) fibrations have $\delta_{\text{tuned}}\geq 200$.
@@
We have plotted the relative density of each of the thirteen fiber types of elliptic fibrations in Figure 11; from this figure, it is clear that fibrations by fiber $F_{13}$ dominate at large $\delta_{r}$. The largest rank increase is $\delta_{r}=463$, found in an $F_{13}$ fibration over base 680 in polytope v05-393, the unique polytope with $h^{1,1}(X)=491$; we discuss this polytope and its fibrations in §5.6. Indeed, this fiber-base pair is universal at large $\delta_{r}$; the largest rank increase that does not come from an $F_{13}$ fibration over base 680 is 350, found in an $F_{13}$ fibration over base 679 in a polytope v06-2907, which has $h^{1,1}(X)=376$, $h^{2,1}(X)=10$. By Eq. 15c, given an elliptic fibration in a CY3 with Hodge numbers $h^{1,1},h^{2,1}$, we can sometimes trade one unit of $h^{1,1}$ for 29 units of $h^{2,1}$. The prevalence of fibrations of $F_{13}$ over base 680 at large $\delta_{r}$ can (at least in part) be explained by the mirror of this process (which will also be described further in Oehlmann and Taylor ): their Hodge numbers are given by $h^{1,1}(X)=491-29n$, $h^{2,1}(X)=11+n$, as expected for the images of the $F_{13}$ fibration in the $h^{1,1}=491$ polytope under the mirror of these $\Delta{T}=1$ transitions. Indeed, in Figure 6 it is clear that fibrations over base 680 dominate the large $h^{1,1}(X)$ regime.
Finally, we note a curious fact: each polytope hosting one of these large rank-increase fibrations of fiber $F_{13}$ over base 680 also possesses a second fibration, namely the generic elliptic fibration over a base with very large $h^{1,1}_{*}(B)$. For instance, as discussed in §5.6, the polytope with $h^{1,1}(X)=491$ also hosts the generic elliptic fibration over the largest toric base. Although we currently cannot explain this empirical phenomenon, we hope that the ideas of Anderson et al. (2016a) can be applied to these pairs of fibrations.